# Gravitational waves from the late inspiral, transition, and plunge of small-mass-ratio eccentric binaries

### v1

@@ Abstract
Black hole binaries with small mass ratios will be important sources for the forthcoming Laser Interferometer Space Antenna (LISA) mission. Models of such binaries also serve as useful tools for understanding the dynamics of compact binary systems and the gravitational waves they emit. Using an eccentric Ori-Thorne procedure developed in previous work, we build worldlines that describe the full inspiral and plunge of a small body on an initially eccentric orbit of a Kerr black hole. We now calculate the gravitational waves associated with these trajectories using a code that solves the Teukolsky equation in the time domain. The final cycles of these waveforms, the ringdown, contains a superposition of Kerr quasinormal modes followed by a power-law tail. In this paper, we study how a binary’s eccentricity and orbital anomaly angle affect the excitation of both quasinormal modes and late-time tails. We find that the relative excitation of quasinormal modes varies in an important and interesting way with these parameters. For some anomaly angles, the relative excitations of quasinormal modes are essentially indistinguishable from those excited in quasi-circular coalescences. Consistent with other recent studies, we find that eccentricity tends to amplify the late-time power-law tail, though the amount of this amplification varies significantly with orbital anomaly. We thus find that eccentricity has an important impact on the late-time coalescence waveform, but the interplay of eccentricity and orbit anomaly complicates this impact.
@@ I.1 Background, motivation, and past work
A merging binary’s coalescence geometry provides critical information about the formation channel which produced the system, a major point of interest for comparable mass systems [19, 20, 21, 22] and small-mass-ratio systems [23] alike. Binaries with significant eccentricity are thought to originate from many-body interactions, such as dynamical encounters in dense stellar environments [24, 25, 26, 27, 28, 29] and three-body interactions [30, 31, 32, 33, 34]. In contrast, aligned-spin systems with small or negligible eccentricity likely form in isolated stellar environments [35, 36, 37]. A diagnostic of eccentricity in the late waveform may serve to enhance the power of GW observations to constrain the relative importance of these different channels.
@@ II.1 The inspiral, transition, and plunge worldline: How it is computed and key parameters describing it
$r_{\rm PLR}=2M\left[1+\cos\left(\frac{2}{3}\cos^{-1}\left[-\frac{|a|}{M}\right]\right)\right]\;.$ (2)
@@ III.1 Quasinormal Modes
As a consequence of the Teukolsky equation’s symmetry, QNMs from Kerr black holes come in pairs: for a given set $(k,m;a,M)$, the eigenvalue problem admits two solutions: $_{s}S_{kmn}^{a\sigma}(\theta,\phi)$, with a positive oscillation frequency, and $_{s}S_{k-mn}^{a\sigma}(\pi-\theta,\phi)^{*}$, with a negative oscillation frequency and a different damping time (where represents complex conjugation). The latter contribution are called “mirror modes” (see [79, 155]). Taking into account contributions from both regular modes and their mirrors, and specializing to spin weight 2-2, we follow [79, 93, 94] and write the QNM contribution to the waveform as
@@ III.1.1 Spherical and spheroidal expansions
To fold mode mixing into our waveform analysis, we first equate the left-hand side of Eq. (7), our numerical data, to the right-hand side of Eq. (10), the QNM contribution to the ringdown. We then multiply both sides of the resultant equation by $_{-2}Y_{\ell m}^{*}(\theta,\phi)$, integrate over the entire sphere, and apply Eq. (11). We then find
@@ III.2 Power-law tails
At sufficiently late times, the ringdown is composed of slowly decaying, power-law controlled tails. In terms of the Teukolsky function $\psi_{4}$, the tail contribution to the ringdown can be modeled as
@@
where $\psi_{4}$ is projected onto the same basis of spherical harmonics we use to describe $h_{\ell m}$. The parameter $\mathcal{A}_{\psi\textrm{tail}}^{\ell m}$ sets the amplitude of the tail when it begins dominating the waveform, $c_{\textrm{tail}}^{\ell m}$ is a time-shift parameter, and $p_{\psi\textrm{tail}}^{\ell m}$ is the power-law index. We expect a power-law index of $p_{\psi\rm tail}^{\ell m}=-(\ell+4)$ [113, 177] to describe the decay of $\psi_{4}]_{\ell m}$. This differs from the power-law index of tails in $h_{\ell m}$, for which we expect $p_{h\rm tail}^{\ell m}=-(\ell+2)$. As we explain in more detail below, we extract tails from $\psi_{4}]_{\ell m}$ rather than $h_{\ell m}$ to eliminate numerical noise effects which can mask tails in our data. We focus on the $\ell=m=2$ harmonic mode, for which we expect $p_{\psi\rm tail}^{22}=-6$. Tails from higher-order spherical modes decay more rapidly, making them difficult to isolate in numerical data. Such contributions, as well as tails from $\ell=2$ modes with $m\neq 2$, are studied at length in IFK25.
@@ IV.2.1 Radial anomaly angle
As we increase the primary’s spin, we continue to see that eccentricity modulates QNM amplitudes depending on the value of $\chi_{r\rm IT}$, though the detailed form is comparatively less “clean” than the Schwarzschild limit. For $a=0.7M$ (bottom row of Fig. 3), the modes oscillate somewhat as $\chi_{r\rm IT}$ increases. The $(2,2)$ mode shows moderately complicated behavior, passing through local extrema as anomaly increases. The general trend, however, is that in eccentric systems, QNM excitation depends in a non-trivial manner on both the eccentricity at the end of inspiral and on the system’s anomaly angle at this moment. For some configurations, mode excitation is similar to the circular limit, with $(2,2)$ dominating. For some anomaly angles, the $(2,1)$ mode dominates, and all modes differ substantially from those found in quasi-circular inspiral.
@@
The worldlines in Fig. 4 reveal a bifurcation in $\chi_{r\rm IT}$ that shapes the late-time dynamics. This “chifurcation” separates two qualitatively distinct kinematic behaviors: systems whose final kinematics have a quasi-circular character before plunge (cf. the solid red curve in Fig. 4, which sits at $r\simeq 4.5M$ for a time interval of several hundred $M$ before plunging), and systems whose final kinematics take the form of a nearly radial infall from relatively large radius (cf. the dot-dashed blue curve in this figure). These distinct behaviors are reminiscent of critical behavior near the unstable circular orbit (UCO) explored in [46, 178]. In particular, the bifurcation appears to be homoclinic [179, 180], as every trajectory seemingly orbits the UCO, then either plunges immediately or completes an additional radial cycle before plunging. Appendix A of Ref. [46] includes a detailed description of homoclinic orbits of Kerr black holes, highlighting the dependence of these orbits on the relativistic anomaly angle at the LSO. In our Appendix B, we study this behavior in detail, including an exploration of how sharply the “chifurcation” depends on mass ratio.
@@ IV.2.2 Mode excitation as a function of the radial velocity at the prograde light ring
Begin with the $a=0$ systems (top row of Fig. 5). The amplitude of the $(2,2)$ mode monotonically declines as $|dr/d\tau|_{r=r_{\rm PLR}}$ grows; all other modes grow with $|dr/d\tau|_{r=r_{\rm PLR}}$. We see the same pattern in systems with $a=0.3M$: as the radial plunge velocity increases, the $(2,2)$ mode weakens while the other $k=2$ mode amplitudes grow with $|dr/d\tau|_{r=r_{\rm PLR}}$. As shown in Appendix A, we tend to find the same behavior among the $k=3$ QNMs: the amplitude of the $(3,3)$ mode falls for high plunge velocities, while other $k=3$ modes grow.
@@ IV.3.2 Tails as a function of the radial anomaly angle
We next examine power-law tails from a set of eccentric binaries, holding all parameters fixed except the late-time radial anomaly angle $\chi_{r\rm IT}$. Both panels of Fig. 8, describe systems with $a=0$, $e_{\rm IT}=0.727$, and $\eta=10^{-4}$. The left-hand panel shows the post-merger amplitude $\psi_{4}]_{22}$ as a function of time after the PLR crossing sourced by the worldlines shown in the right-hand panel. These binaries are distinct in both in their late-time behavior and their excited ringdown amplitudes, despite the fact that they only differ by $\chi_{r\rm IT}$. Table 2 gives the tail parameters we find for these systems. Visual inspection of the tails paired with the fit parameters we find shows that a binary’s terminal dynamics strongly shapes the morphology of its power-law tail.
@@ V Conclusions and Next Steps
Secondaries that plunge after completing quasi-circular whirls excite QNM spectra that closely mimic the mode structure of purely quasi-circular mergers. In particular, the $(2,2)$ mode dominates. These systems can have significant eccentricity during the inspiral, but if inspiral ends at an anomaly angle such that the system whirls at periapse before plunging, the ringdown appears quasi-circular. If the same secondary approaches the strong field with a different anomaly, it may instead plunge from large radius after a final radial cycle. In this case, the spectrum is dominated by the $(2,1)$ QNM. In addition to depending on anomaly angle, these behaviors depend on the system’s spin. High spins make $(2,1)$ dominance possible at lower eccentricities.
@@ Appendix A Higher Order Mode Excitation
In Sec. IV, we explored the excitation of QNMs and late-time tails as a function of the radial anomaly at the end of adiabatic inspiral, $\chi_{r\rm IT}$. Ringdowns from eccentric systems that differ only by $\chi_{r\rm IT}$ can present distinct QNM amplitudes and tail morphologies. Further, in Sec. IV.2.1 we found a region of parameter space in which the worldline is especially sensitive to $\chi_{r\rm IT}$, resulting in a bifurcation in the late-time kinematics that is particularly consequential to the ringdown. This bifurcation with respect to $\chi_{r\rm IT}$, which we have dubbed the “chifurcation,” represents an extreme case of an eccentric inspiral’s sensitivity to initial conditions. In this appendix, we study the chifurcation in more detail, focusing on its dependence on the binary’s mass ratio.

### v2

@@ Abstract
Black hole binaries with small mass ratios will be important sources for the forthcoming Laser Interferometer Space Antenna (LISA) mission. Models of such binaries also serve as useful tools for understanding the dynamics of compact binary systems and the gravitational waves they emit. Using an eccentric Ori-Thorne procedure developed in previous work, we build worldlines that describe the full inspiral and plunge of a small body on an initially eccentric orbit of a Kerr black hole. We now calculate the gravitational waves associated with these trajectories using a code that solves the Teukolsky equation in the time domain. The final cycles of these waveforms, the ringdown, contain a superposition of Kerr quasinormal modes followed by a power-law tail. In this paper, we study how a binary’s eccentricity and orbital anomaly angle affect the excitation of both quasinormal modes and late-time tails. We find that the relative excitation of quasinormal modes varies in an important and interesting way with these parameters. For some anomaly angles, the relative excitations of quasinormal modes are essentially indistinguishable from those excited in quasi-circular coalescences. Consistent with other recent studies, we find that eccentricity tends to amplify the late-time power-law tail, though the amount of this amplification varies significantly with orbital anomaly. We thus find that eccentricity has an important impact on the late-time coalescence waveform, but the interplay of eccentricity and orbit anomaly complicates this impact.
@@ I.1 Background, motivation, and past work
A merging binary’s coalescence geometry provides critical information about the formation channel which produced the system, a major point of interest for comparable mass systems [19, 20, 21, 22] and small-mass-ratio systems [23] alike. Binaries with significant eccentricity are thought to originate from many-body interactions, such as dynamical encounters in dense stellar environments [24, 25, 26, 27, 28, 29] and three-body interactions [30, 31, 32, 33, 34, 35, 36, 37]. In contrast, aligned-spin systems with small or negligible eccentricity likely form in isolated stellar environments [38, 39, 40]. A diagnostic of eccentricity in the late waveform may serve to enhance the power of GW observations to constrain the relative importance of these different channels.
@@ II.1 The inspiral, transition, and plunge worldline: How it is computed and key parameters describing it
$r_{\rm PLR}=2M\left[1+\cos\left(\frac{2}{3}\cos^{-1}\left[-\frac{a}{M}\right]\right)\right]\;.$ (2)
@@ III.1 Quasinormal Modes
As a consequence of the Teukolsky equation’s symmetry, QNMs from Kerr black holes come in pairs: for a given set $(k,m;a,M)$, the eigenvalue problem admits two solutions: $_{s}S_{kmn}^{a\sigma}(\theta,\phi)$, with a positive oscillation frequency, and $_{s}S_{k-mn}^{a\sigma}(\pi-\theta,\phi)^{*}$, with a negative oscillation frequency and a different damping time (where represents complex conjugation). The latter contributions are called “mirror modes” (see [82, 158]). Taking into account contributions from both regular modes and their mirrors, and specializing to spin weight 2-2, we follow [82, 96, 97] and write the QNM contribution to the waveform as
@@ III.1.1 Spherical and spheroidal expansions
To fold mode mixing into our waveform analysis, we first equate the left-hand side of Eq. (7), our numerical data, to the right-hand side of Eq. (10), the QNM contribution to the ringdown. We then multiply both sides of the resultant equation by $_{-2}Y_{\ell m}^{*}(\theta,\phi)$, integrate over the sphere, and apply Eq. (11). We then find
@@ III.2 Power-law tails
At sufficiently late times, the ringdown is composed of slowly decaying, power-law tails. In terms of the Teukolsky function $\psi_{4}$, the tail contribution to the ringdown can be modeled as
@@
where $\psi_{4}$ is projected onto the same basis of spherical harmonics we use to describe $h_{\ell m}$. The parameter $\mathcal{A}_{\psi\textrm{tail}}^{\ell m}$ sets the amplitude of the tail when it begins dominating the waveform, $c_{\textrm{tail}}^{\ell m}$ is a time-shift parameter, and $p_{\psi\textrm{tail}}^{\ell m}$ is the power-law index. We expect a power-law index of $p_{\psi\rm tail}^{\ell m}=-(\ell+4)$ [116, 180] to describe the decay of $\psi_{4}]_{\ell m}$. This differs from the power-law index of tails in $h_{\ell m}$, for which we expect $p_{h\rm tail}^{\ell m}=-(\ell+2)$. As we explain in more detail below, we extract tails from $\psi_{4}]_{\ell m}$ rather than $h_{\ell m}$ to avoid numerical noise effects which can mask tails in our data. We focus on the $\ell=m=2$ harmonic mode, for which we expect $p_{\psi\rm tail}^{22}=-6$. Tails from higher-order spherical modes decay more rapidly, making them difficult to isolate in numerical data. Such contributions, as well as tails from $\ell=2$ modes with $m\neq 2$, are studied at length in IFK25.
@@ IV.2.1 Radial anomaly angle
As we increase the primary’s spin, we continue to see that eccentricity modulates QNM amplitudes depending on the value of $\chi_{r\rm IT}$, though the detailed form is comparatively less “clean” than in the Schwarzschild limit. For $a=0.7M$ (bottom row of Fig. 3), the modes oscillate somewhat as $\chi_{r\rm IT}$ increases. The $(2,2)$ mode shows moderately complicated behavior, passing through local extrema as anomaly increases. The general trend, however, is that in eccentric systems, QNM excitation depends in a non-trivial manner on both the eccentricity at the end of inspiral and on the system’s anomaly angle at this moment. For some configurations, mode excitation is similar to the circular limit, with $(2,2)$ dominating. For some anomaly angles, the $(2,1)$ mode dominates, and all modes differ substantially from those found in quasi-circular inspiral.
@@
The worldlines in Fig. 4 reveal a bifurcation in $\chi_{r\rm IT}$ that shapes the late-time dynamics. This “chifurcation” separates two qualitatively distinct kinematic behaviors: systems whose final kinematics have a quasi-circular character before plunge (cf. the solid red curve in Fig. 4, which sits at $r\simeq 4.5M$ for a time interval of several hundred $M$ before plunging), and systems whose final kinematics take the form of a nearly radial infall from relatively large radius (cf. the dot-dashed blue curve in this figure). These distinct behaviors are reminiscent of critical behavior near the unstable circular orbit (UCO) explored in [49, 181]. In particular, the bifurcation appears to be homoclinic [182, 183], as every trajectory seemingly orbits the UCO, then either plunges immediately or completes an additional radial cycle before plunging. Appendix A of Ref. [49] includes a detailed description of homoclinic orbits of Kerr black holes, highlighting the dependence of these orbits on the relativistic anomaly angle at the LSO. In both our work and Ref. [49], the chifurcation for a nonspinning system appears to occur near the last periapsis passage, which is consistent with the character of periapsis, since this is where the secondary switches from moving radially toward the primary to moving away from it. We can also schematically understand this approximate chifurcation location with an effective potential description of the motion (see Ref. [49] for a more detailed discussion of strong-field black hole orbits in effective potentials and their sensitivity to the anomaly angle). In Appendix B, we investigate this behavior in detail with an exploration of how sharply the “chifurcation” depends on mass ratio, and a brief study of the influence of spin and eccentricity at plunge on the location of the chifurcation.
@@ IV.2.2 Mode excitation as a function of the radial velocity at the prograde light ring
Begin with the $a=0$ systems (top row of Fig. 5). The amplitude of the $(2,2)$ mode monotonically declines as $|dr/d\tau|_{r=r_{\rm PLR}}$ grows; all other modes grow with $|dr/d\tau|_{r=r_{\rm PLR}}$. We see the same pattern in systems with $a=0.3M$: as the radial plunge velocity increases, the $(2,2)$ mode weakens while the other $k=2$ mode amplitudes grow with $|dr/d\tau|_{r=r_{\rm PLR}}$. As shown in Appendix A, we tend to find analogous behavior among the $k=3$ QNMs: the amplitude of the $(3,3)$ mode falls for high plunge velocities, while other $k=3$ modes grow.
@@ IV.3.2 Tails as a function of the radial anomaly angle
We next examine power-law tails from a set of eccentric binaries, holding all parameters fixed except the late-time radial anomaly angle $\chi_{r\rm IT}$. Both panels of Fig. 8, describe systems with $a=0$, $e_{\rm IT}=0.727$, and $\eta=10^{-4}$. The left-hand panel shows the post-merger amplitude $\psi_{4}]_{22}$ as a function of time after the PLR crossing sourced by the worldlines shown in the right-hand panel. These binaries are distinct both in their late-time behavior and their excited ringdown amplitudes, despite the fact that they only differ by $\chi_{r\rm IT}$. Table 2 gives the tail parameters we find for these systems. Visual inspection of the tails paired with the fit parameters we find shows that a binary’s terminal dynamics strongly shapes the morphology of its power-law tail.
@@ V Conclusions and Next Steps
Secondaries that plunge immediately after completing quasi-circular whirls excite QNM spectra that closely mimic the mode structure of purely quasi-circular mergers. In particular, the $(2,2)$ mode dominates. These systems can have significant eccentricity during the inspiral, but if inspiral ends at an anomaly angle such that the system whirls at periapse before plunging, the ringdown appears quasi-circular. If the same secondary approaches the strong field with a different anomaly, it may instead plunge from large radius after a final radial cycle. In this case, the spectrum is dominated by the $(2,1)$ QNM. In addition to depending on anomaly angle, these behaviors depend on the system’s spin. High spins make $(2,1)$ dominance possible at lower eccentricities.
@@ Appendix A Higher Order Mode Excitation
In Sec. IV, we explored the excitation of QNMs and late-time tails as a function of the radial anomaly at the end of adiabatic inspiral, $\chi_{r\rm IT}$. Ringdowns from eccentric systems that differ only by $\chi_{r\rm IT}$ can present distinct QNM amplitudes and tail morphologies. Further, in Sec. IV.2.1 we found a region of parameter space in which the worldline is especially sensitive to $\chi_{r\rm IT}$, resulting in a bifurcation in the late-time kinematics that is particularly consequential to the ringdown. This bifurcation with respect to $\chi_{r\rm IT}$, which we have dubbed the “chifurcation,” represents an extreme case of an eccentric inspiral’s sensitivity to initial conditions. In this appendix, we study the chifurcation in more detail, first focusing on its dependence on the binary’s mass ratio.
@@
We also carried out a brief investigation into the influence of a system’s mass ratio, spin and late-time eccentricity on the chifurcation angle. Although we always find that the chifurcation occurs slightly before periapsis, its precise location can differ between systems with different $\eta,a$ and $e_{\rm IT}$. In Table 3, we show angular intervals $X^{\rm chif}_{r\rm{IT}}$ that bracket the chifurcation for a variety of systems. For a given $\eta,a$ and $e_{\rm IT}$, systems with $\chi_{r\rm{IT}}$ at the lower bound of $X^{\rm crit}_{r\rm{IT}}$ execute quasi-circular whirls before plunging, while systems with $\chi_{r\rm{IT}}$ at the upper bound of $X^{\rm crit}_{r\rm{IT}}$ move radially outward before plunging. The transition between these two behaviors, the chifurcation, thus occurs at some $\chi_{r\rm{IT}}\in X^{\rm crit}_{r\rm{IT}}$. While narrowing the width of each interval $X^{\rm chif}_{r\rm{IT}}$ is possible with repeated calculations, the intervals do not overlap for the cases we wish to compare, so the given resolution of $X^{\rm chif}_{r\rm{IT}}$ is sufficient. The $X^{\rm chif}_{r\rm{IT}}$ intervals in Table 3 indicate that increasing the spin shifts the chifurcation to smaller $\chi_{r\rm IT}$, while increasing the eccentricity shifts the chifurcation to larger $\chi_{r\rm IT}$. Although we only examined two mass ratios, $\eta=10^{-4}$ and $\eta=3\times 10^{-4}$, what we see is that, holding all other parameters constant, making the mass ratio less extreme tends to move the chifurcation to a point earlier than periapsis (i.e., to a somewhat smaller value of $\chi^{\rm chif}_{r\rm{IT}}$).
$\eta$ $a/M$ $e_{\rm IT}$   Bracketing interval $X^{\textrm{chif}}_{r\textrm{IT}}$ [deg]
  $10^{-4}$   0   0.27    $353.18<\chi_{r\rm IT}^{\rm chif}<353.34$
$10^{-4}$   0   0.46    $355.24<\chi_{r\rm IT}^{\rm chif}<355.40$
$10^{-4}$   0   0.67    $356.27<\chi_{r\rm IT}^{\rm chif}<356.34$
  $10^{-4}$   0.1   0.27    $352.80<\chi_{r\rm IT}^{\rm chif}<352.96$
$10^{-4}$   0.1   0.46    $354.98<\chi_{r\rm IT}^{\rm chif}<355.11$
$10^{-4}$   0.1   0.67   $356.08<\chi_{r\rm IT}^{\rm chif}<356.18$
  $10^{-4}$   0.3   0.27    $352.29<\chi_{r\rm IT}^{\rm chif}<352.48$
$10^{-4}$   0.3   0.46    $354.72<\chi_{r\rm IT}^{\rm chif}<354.87$
$10^{-4}$   0.3   0.67    $355.85<\chi_{r\rm IT}^{\rm chif}<355.95$
  $10^{-4}$   0.7   0.27    $350.71<\chi_{r\rm IT}^{\rm chif}<350.93$
$10^{-4}$   0.7  0.46    $353.81<\chi_{r\rm IT}^{\rm chif}<353.96$
$10^{-4}$   0.7   0.67    $355.21<\chi_{r\rm IT}^{\rm chif}<355.30$
  $3\times 10^{-4}$   0   0.27    $350.32<\chi_{r\rm IT}^{\rm chif}<350.59$
$3\times 10^{-4}$   0  0.67    $354.80<\chi_{r\rm IT}^{\rm chif}<355.01$
  $3\times 10^{-4}$   0.3  0.27    $349.28<\chi_{r\rm IT}^{\rm chif}<349.55$
$3\times 10^{-4}$   0.3   0.67    $354.16<\chi_{r\rm IT}^{\rm chif}<354.37$
  $3\times 10^{-4}$   0.7  0.27    $347.06<\chi_{r\rm IT}^{\rm chif}<347.36$
$3\times 10^{-4}$   0.7  0.67    $353.19<\chi_{r\rm IT}^{\rm chif}<353.40$
Table 3: Angular intervals $X^{\rm chif}_{r\rm{IT}}$ that bracket the chifurcation for systems with different $\eta,a$ and $e_{\rm IT}$.